OpenStudy (anonymous):

explain what the discriminant is and what it tells you about the solutions of a quadratic equation

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

the discriminant is b^2 - 4ac it tells u how many roots/solutions an equation has and if they are real or imaginary

OpenStudy (paxpolaris):

quadratic eqauation: \(\Large ax^2+bx+c=0\) \[\Large x= {-b \pm \sqrt{b^2-4ac} \over 2a}\] Discriminant: \(\Large D=b^2-4ac\)

OpenStudy (paxpolaris):

when D>0 ... you have 2 distinct real number solutions when D=0 ... you have 1 repeated real solution when D<0 ... you have 0 real number solutions ... or 2 complex number solution as long as a,b,c are real numbers, and a is not 0